OpenStudy (anonymous):

N+ 3/4 = 20 can you tell me the value of N? thank u

OpenStudy (nottim):

It's algebra.

OpenStudy (nottim):

Do the orders of BEDMAS, or PEDMAS, in reverse.

OpenStudy (nottim):

Think this; how would you isolate "n"?

OpenStudy (anonymous):

4N+3=80 4N=77 N=77/4

jhonyy9 (jhonyy9):

N=20 - 3/4 = 77/4

OpenStudy (nottim):

Jhonny, you keep doing this!!

OpenStudy (anonymous):

i am an idiot.. sorry.. it is n +3/4 = 19/20

OpenStudy (nottim):

erps

OpenStudy (anonymous):

i cant do fractions and I cant type// sorry

OpenStudy (nottim):

But were you able to learn from the prev. examples then?

OpenStudy (nottim):

The process is the same.

OpenStudy (anonymous):

my answer is 1/2 but not sure if im correct?

OpenStudy (nottim):

lemme try it out...

OpenStudy (nottim):

I got 0.2...

OpenStudy (nottim):

Mind showing me what you did?

OpenStudy (anonymous):

can you type the problem again

OpenStudy (anonymous):

N= 19/20 - 3/4 19/20 - 15/20 4/20 = 2/10 = 1/2

OpenStudy (nottim):

2/10 does not factor down to 1/2

OpenStudy (anonymous):

The problem is: N + 3/4 = 19/20

OpenStudy (nottim):

But earlier, you did it correctly.

OpenStudy (nottim):

2/10 does factor down to 1/5, as both 2 and 10 are divided by 2.

OpenStudy (anonymous):

ask this website called tutor.com

OpenStudy (nottim):

do you have to pay for it?

OpenStudy (anonymous):

they really help

OpenStudy (anonymous):

you dont have to pay

OpenStudy (nottim):

awesome. Imma use that too.

OpenStudy (anonymous):

but you have to sigh up

OpenStudy (nottim):

So, lamby, you know how to do it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i will.. can u please tell me if my steps until 2/10 are correct? or am I completely wrong?

OpenStudy (nottim):

All steps are correctly, up to 2/10.

OpenStudy (anonymous):

but tutor .com is only open at 2:00

OpenStudy (nottim):

Aside from your error in simplification in the final steps, you did it correctly.

OpenStudy (anonymous):

ok.. thank you

OpenStudy (nottim):

Remember, what you do to the top is the same thing you do to the bottom.

OpenStudy (anonymous):

your welcome

OpenStudy (nottim):

Can I get a medal? That all my life is worth now.

OpenStudy (nottim):

Thanks.

OpenStudy (anonymous):

i want one

OpenStudy (anonymous):

2

OpenStudy (anonymous):

i did....

OpenStudy (nottim):

Same here.

OpenStudy (anonymous):

i gave medals to all of you for yor time and help.. thank u